OpenStudy (anonymous):

Given f(x)=17-x^2 what is the average rate of change in f(x) over the interval [1, 5]?

OpenStudy (amistre64):

the slope of the line that connects the endpoints ....

OpenStudy (anonymous):

What do you mean by that? i dont understand

OpenStudy (amistre64):

you find the slope of the line that connects the ends points

OpenStudy (amistre64):

given 2 points, how do you find the slope of a line between them?

OpenStudy (anonymous):

Don,t you count the sqaures?

OpenStudy (anonymous):

Squares*

OpenStudy (amistre64):

hmm, thats a graphical approach yes. which may or maynot be applicable here

OpenStudy (anonymous):

There is no graph involved

OpenStudy (amistre64):

if we go by the slope formula: lets find the change in y as it relates to the change in x\[slope:~\frac{y_1-y_o}{x_1-x_o}\]

OpenStudy (amistre64):

in this case, we have our y parts as f(a) and f(b) for an interval from a to b

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

writing this down

OpenStudy (amistre64):

with any luck it already written down in your course material

OpenStudy (anonymous):

Well of course, but i never quite understood it. im trying to find an easier way to understand, which is why i came here.

OpenStudy (amistre64):

there is no easier way to understand it. practice is the only way to make it understable

OpenStudy (anonymous):

Okay, so in the equation would 17 be x? to find the slope..

OpenStudy (amistre64):

we know x1 and x2: 1 and 5, the endpoints of our interval we want to know f(1) and f(5), which will be our y values to compare

OpenStudy (amistre64):

the 17 part here is superfluous, but that might just confuse you at the moment

OpenStudy (anonymous):

wait, are we subtracting?

OpenStudy (amistre64):

yes, that will be the only way to find the difference between things, subtract them

OpenStudy (amistre64):

\[slope:~~\frac{f(1)-f(5)}{1-5}\]

OpenStudy (anonymous):

OH! the answer is 1?!

OpenStudy (amistre64):

i doubt it

OpenStudy (anonymous):

ughhh! i thought i had it.

OpenStudy (amistre64):

we need to know the value of f(1) and f(5)

OpenStudy (anonymous):

because i got negative 4 on top and bottom then i divided them and got 1

OpenStudy (anonymous):

Okay

OpenStudy (amistre64):

let me dbl chk f(1) = 17-(1)^2 f(5) = 17-(5)^2 f(1) - f(5) is: 17-(1)^2 -(17-(5)^2) = -1+25

OpenStudy (amistre64):

17-17 is 0 which is why a foretold that the 17 is really useless

OpenStudy (amistre64):

so \[\frac{f(1)-f(5)}{1-5}=\frac{-1+25}{-4}=\frac{24}{-4}\]

OpenStudy (anonymous):

how did you get -1+25 on the top??

OpenStudy (amistre64):

do you know what to do with a function? the f(x) that was defined?

OpenStudy (anonymous):

I'm horrible at functions so no

OpenStudy (amistre64):

well its really quite simple; lets name the equation as f(x), read 'f of x'. lets give the name a definition, a rule, an equation f(x) = 17 - x^2 now, for any value of x, we can determine the value of f(x) let x=1 f(1) = 17-1^2 let x=5 f(5) = 17 - 5^2

OpenStudy (amistre64):

f(1) - f(5) = (17-1)-(17-25) = -1+25